[Hallc_running] [New Logentry] Hall C Status Update

latif at jlab.org latif at jlab.org
Mon Jan 22 13:05:01 EST 2018

Logentry Text:
The target people went in and found the target frame has melted. They will be removing the frame.
Based on the current situation,  today's swing shift and owl shift  are cancelled. The day shift tomorrow will remain on standby. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3519135
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180122/4879be42/attachment.html>

More information about the Hallc_running mailing list